At noon, two trains leave from towns 175 miles apart. One train averages 40 mph and the other averages 30 mph. At what time do they meet? (Remeber: distance= rate*time)
Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

40X+30X=175
70X=175
X=175/70
X=2.5 OR 2:30 PM THEY WILL MEET.

Answer by stanbon(75887)

At noon, two trains leave from towns 175 miles apart. One train averages 40 mph and the other averages 30 mph. At what time do they meet? (Remeber: distance= rate*time)
----------------
Train One DATA:
Distance: x : Rate= 40 mph ; Time = d/r = x/40 hrs
========
Train Two DATA:
Distance: 175-x miles ; Rate = 30 mph ; Time = d/r = (175-x)/30 hrs
----------------
EQUATION:
Time DISABLED_event_One= Time Two
x/40 = (175-x)/30
Cross-multiply to get:
30x = 40(175-x)
3x = 4(175-x)
3x = 700 -4x
7x = 700
x=100 miles
Time = x/40 = 100/40 = 2.5 hrs
===============
Cheers,
Stan H.